OpenStudy (anonymous):

integrate from 0 to 3. x e^x^2 dx

OpenStudy (anonymous):

i have the integration = 1/2 e^x^2 + c

OpenStudy (anonymous):

e^(9)/2-(1/2)

OpenStudy (anonymous):

That's right, now evaluate it using your limits.

OpenStudy (anonymous):

thats the thing

OpenStudy (anonymous):

what happens when e goes to 9

OpenStudy (anonymous):

i know that when e goes to 0 gos to negative infinitive

myininaya (myininaya):

e^0=1

OpenStudy (anonymous):

\(e^9\) is just a number, like \(\pi\) etc. \(e^0 = 1\) Just like any other base raised to the 0 power equals 1.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

so e ^9 is postive infinitive???

OpenStudy (anonymous):

No, it's just a number.

myininaya (myininaya):

no

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

e*e*e*e*e*e*e*e*e

OpenStudy (anonymous):

maybe i am confusing this problem with the limits

myininaya (myininaya):

e is approximately 2.7 so e^9 is approximately (2.7)^9

OpenStudy (anonymous):

ohhhh ok

myininaya (myininaya):

just like pi is just a number

OpenStudy (anonymous):

And \(\sqrt{2}\)

OpenStudy (anonymous):

cool,,